A COFFEE MERCHANT COMNBINED COFFEE COSTING $6 PER POUND, WITH COFFEE COSTING $3.50 PER POUND. HOW MANY POUNDS OF EACH WERE USED TO MAKE 25 POUNDS OF A BLEND COSTING $5.25 PER POUND?

Answer by checkley77(12844)

6x+3.5(25-x)=25*5.25
6x+87.5-3,5x=131.25
2.5x=131.25-87.5
2.5x=43.75
x=43.75/2.5
x=17.5 lbs. of $6 coffee is used.
25-17.5=7.5 lbs. of $3.50 coffee is used.
Proof:
6*17.5+3.5*7.5=131.25
105+26.25=131.25
131.25=131.25